Two seventh of the people in a park are children. There are 45 more adults than children. How many people are there altogether?
Answer by Maths68(1474)

Let
Children = c
Adult = a
Two seventh of the people in a park are children
(Adult+children)(2/7)=children
(a+c)(2/7)=c
(2)(a+c)/7=c
(2)(a+c)=c*7.................(1)
There are 45 more adults than children
a=c+45.......................(2)
Substitute the value of a from (2) to (1)
(2)(a+c)=c*7.................(1)
(2)(c+45+c)=c*7
2c+90+2c=7c
4c+90=7c
4c-7c=-90
-3c=-90
-3c/-3=-90/-3
c=30
Put the value of c in (2)
a=c+45.......................(2)
a=30+45
a=75
Children = c = 30
Adult = a = 75